Roopa Patel won a prize for her sculpture. She invested part of the money in a certificate of deposit at 4% and $3000 more than that amount in a bond paying 6%. Her annual interest income was $780. How much did Roopa invest at each rate?
Answer by mananth(16946)

Roopa Patel won a prize for her sculpture. She invested part of the money in a certificate of deposit at 4% and $3000 more than that amount in a bond paying 6%. Her annual interest income was $780. How much did Roopa invest at each rate?
...
$ x in CD @ 4%
$(x+3000) in bond @ 6 %
total interest income in 1 year was $780
...
Interest equation
0.04x+0.06x(x+3000)= 780
0.04x+0.06x+180 = 780
0.1x = 780-180
0.1x=600
x=600/0.1
x= $6000 @ 4%
$3000 more @6%
6000+3000=$9000 @6%
...
m.ananth@hotmail.ca